multiply and simplfy: √8*√54
Answer by dabanfield(803)

multiply and simplfy: √8*√54
sqrt(8)*sqrt(54) = sqrt(2*2*2)*sqrt(3*3*3*2) =
sqrt(2*2*2*3*3*3*2) = sqrt(2*2)*sqrt(2*2)*sqrt(3*3)*sqrt(3) =
2*2*3*sqrt(3) = 12*sqrt(3)